# An End-to-End Model For Logits Based Large Language Models Watermarking

### v1

@@ Abstract
The rise of LLMs has increased concerns over source tracing and copyright protection for AIGC, highlighting the need for advanced detection technologies. Passive detection methods usually face high false positives, while active watermarking techniques using logits or sampling manipulation offer more effective protection. Existing LLM watermarking methods, though effective on unaltered content, suffer significant performance drops when the text is modified and could introduce biases that degrade LLM performance in downstream tasks. These methods fail to achieve an optimal tradeoff between text quality and robustness, particularly due to the lack of end-to-end optimization of the encoder and decoder. In this paper, we introduce a novel end-to-end logits perturbation method for watermarking LLM-generated text. By jointly optimization, our approach achieves a better balance between quality and robustness. To address non-differentiable operations in the end-to-end training pipeline, we introduce an online prompting technique that leverages the on-the-fly LLM as a differentiable surrogate. Our method achieves superior robustness, outperforming distortion-free methods by 37–39% under paraphrasing and 17.2% on average, while maintaining text quality on par with these distortion-free methods in terms of text perplexity and downstream tasks. Our method can be easily generalized to different LLMs.
@@ 3.3 Online Text Editing
As shown in Fig. 1 (b), the online text editor $N$ (in light purple) is positioned between the encoder and decoder during training to simulate user edits of watermarked content, enhancing detection robustness. We utilize the online prompting technique in Fig. 2 (a) to augment watermarked text by dynamically prompting the online LLM $M$. This approach effectively handles non-differentiable online text modifications, such as rewriting. Specifically, after generates watermarked text $\mathbf{X}_{\text{wm}}$, which is fed into $N$ to produce $\hat{\mathbf{X}}_{\text{wm}}$. The editing prompt: Rewrite the following paragraph: [text] . Now start to rewrite the above paragraph: instructs $M$ to generate an augmented version of the watermarked content, where [text] is a placeholder for $\mathbf{X}_{\text{wm}}$. Thus, $N(\mathbf{X}_{\text{wm}})$ is equivalent to $M([\mathbf{X}_{\text{epb}},\mathbf{X}_{\text{wm}},\mathbf{X}_{\text{epe}}])$ in which $\mathbf{X}_{\text{epb}}$ and $\mathbf{X}_{\text{epe}}$ denote the beginning and the end of the editing prompt. In contrast to existing online text editing methods, such as random token dropping/adding (Zhang et al., 2024), our approach shows significantly improved robustness to unseen distortions due to the capabilities of the on-the-fly $M$ to perform more complex modification on the text (more examples are provided in Appendix K). While external paraphrasing models like Dipper (Krishna et al., 2023) are available, differences in tokenizers between the LLM and these models render the operation non-differentiable.
@@ 3.6 Cross-LLM Inference
Robustness Quality Robustness Quality Robustness Quality
@@ 4.1.1 Implementation Details
To train our end-to-end model, we choose OPT-1.3B as the online LLM for efficiency. We use samples from the WikiText-103 dataset (Merity et al., 2017) as prompts during training and use samples from the C4 (Raffel et al., 2020) dataset during the evaluation. The context window is set to be $W=10$, with the watermark strength $\delta=1.25$ and $k=20$ for the top-$k$ watermark logits empirically. We employ MGDA (Huo et al., 2024; Désidéri, 2012) to balance detection and semantic losses and use the Adam optimizer with a learning rate of 1e-4 over 35k training steps. All experiments are conducted on one single NVIDIA RTX A6000 48G GPU. The evaluation is performed based on the MarkLLM [^2] https://github.com/THU-BPM/MarkLLM (Pan et al., 2024), an open-source tool for benchmarking LLM watermark methods. For further details, please refer to our code in supplementary material, as well as the training details in Appendix F, and the editing configuration in Appendix G.
@@ 4.1.2 Metrics
We benchmark LLM watermarking methods across four key aspects, following the tradition of Pan et al. (2024); Zhang et al. (2024): 1) Detection effectiveness of the human-written and clean watermarked text, measured by the F1 scores at optimal thresholds; 2) Detection robustness, evaluated by subjecting the watermarked text to 3 types of text modifications; 3) Text quality, assessed using the text perplexity as well as performance on downstream tasks including translation and code generation; and 4) Efficiency of the watermarking model, evaluated by measuring the time and GPU memory overhead for text generation and watermark detection, as detailed in Appendix J.
@@ 4.1.3 Competitors
We compare our method with the SOTA logits-based methods: KGW(Kirchenbauer et al., 2023), SIR(Liu et al., 2024b), and Unigram(Zhao et al., 2024), as well as distortion-free methods: Unbiased(Hu et al., 2024) and DiPmark(Wu et al., 2024). A comparison involve more competitors is presented in Appendix A.
@@ 4.2 Robustness Quality Benchmarking
Table 2 compares the robustness and quality of five methods across three LLMs, with Unigram and DiPmark serving as strong baselines. The performance gap ($\Delta$) is presented at the bottom of the table. Using the first 30 tokens from the C4 dataset (Raffel et al., 2020) as prompts, 200 clean watermarked tokens are generated, with original human-written text serving as non-watermarked samples. Robustness is evaluated under clean and edited conditions (SS, CP, PA in Table 2), while text quality is assessed using PPL with Llama2-13B as the oracle. Unlike Unbiased, which relies on prompts and LLM API access, making it inefficient, our method and other baselines operate without such constraints. Notably, our model is trained exclusively on OPT-1.3B and employs the converter (Section 3.6) for cross-LLM inference. Our method outperforms all baselines in robustness and quality across most scenarios, with minor degradations (1–3%) in a few cases, achieving an average F1 score of 0.975 across all scenarios. Our method achieves an average robustness improvement of 17.2% over DiPmark, with notable gains of 37–39% under paraphrasing, while maintaining comparable PPL scores. Compared to Unigram, our approach shows a 1.5% average robustness gain, peaking at 12% under copy-paste attacks for Llama2-7B, and reduces PPL by over 9% on average. These results highlight that our method effectively delivers watermark resilience while preserving LLM output quality, making it a practical solution for real-world applications.
@@
To further validate the LLM-agnostic generalizability of our method, we conduct experiments on three additional LLMs, with the results summarized in Table 3. Our approach consistently achieves a high F1 score ($\geq 0.99$) on clean watermarked samples across all LLMs, demonstrating its stability and reliability. Moreover, it maintains strong robustness against all edited cases, yielding an average F1 score of 0.969. In terms of text quality, our method introduce only a moderate increase in PPL, approximately 1.2× that of non-watermarked baselines. These zero-shot results highlight the effectiveness of our approach without necessitating model-specific tuning, underscoring its broad applicability across diverse LLM architectures.
@@
NLLB-600M
@@
Starcoder
@@
The strength parameter $\delta$ in logits-based methods is a critical factor in balancing robustness and quality. To systematically evaluate this trade-off, we conduct experiments assessing watermark resilience against paraphrasing (PA) and the copy-paste attack while measuring PPL. Watermarked text is generated under varying $\delta$ values [^3] Each method is tested at its default strength, as well as one lower and one higher setting: KGW and Unigram: {1, 2, 5}; SIR: {0.5, 1, 2}; Ours: {0.75, 1.25, 1.5}. The results, presented in Fig. 5, demonstrate the superior performance of our method in achieving an optimal balance between robustness and quality. At a PPL threshold of 8, where the non-watermarked baseline has a PPL of 6.8, our approach improves robustness by 18.16% under paraphrasing and 20.42% under copy-paste edits, surpassing the second-best method, Unigarm. In contrast, existing approaches such as KGW, Unigram, and SIR exhibit significant text quality degradation, with PPL values increasing by two to three times compared to the default $\delta$ while offering only marginal robustness improvements. These results highlight the effectiveness of our method in preserving text quality while achieving substantial robustness against text modifications.
Additional experiments are detailed in the appendix. Appendix A benchmarks more competitors, while detection performance is further evaluated across different sentence lengths in Appendix B, under varying LLM temperatures in Appendix C, and across different sampling strategies in Appendix D.
@@ 4.3 Comparsion to KGW-type Methods
We justify the proposed neural-based methods by comparing them to KGW-type approaches, highlighting that our method offers guarantees analogous to provable p-values, which are essential for text forensics.
Red-Green List Retrieval. Although our method does not explicitly rely on the p-value to determine the presence of the watermark, we allow for forensic analysis by retrieving the logits perturbation, which serves as the counterpart to the red-green list in KGW, for each token in a given sentence. Specifically, given an input sequence $[\mathbf{x}_{1},\dots,\mathbf{x}_{n}]$, tokenization is performed using the proposed converter. The retrieval is achieved by first obtaining the output distribution of $M$ at each step for $[\mathbf{x}_{w},\dots,\mathbf{x}_{n}]$ and subsequently reconstructing $\mathbf{S}_{\text{top}\,k}^{(t)}$ for each step $t$. The logits perturbation are then sequentially derived as $\mathbf{l}^{(t)}_{\text{top}\,k}=E(\mathbf{S}_{\text{top}\,k}^{(t)})$. Our watermarked samples colored with retrieval logits perturbation are visualized in Appendix E, this retrieval process enables the classification of tokens into green and red lists, thereby enhancing the interpretability of our watermark signal.
@@
Statistical Decoder vs. Neural Decoder. Our method enables the retrieval of logits perturbation for a given sentence, allowing us to visualize the relationship between cumulative perturbed logits and neural decoder logits in Fig. 6. The analysis shows a strong positive correlation, where neural decoder logits increase as cumulative perturbed logits rise across the two LLMs. Additionally, the neural decoder outperforms the statistical decoder, likely due to the joint optimization process, which allows it to capture latent features that enhance detection performance.
False Positive Thresholding. Our model uses a neural decoder to generate a scalar confidence score for watermark detection. Similar to KGW-type methods, both employ a fixed threshold to detect watermarks, enabling adjustments to control false positive rates. Fig. 7 shows ROC curves under various text modifications, demonstrating the model’s ability to manage false positive rates effectively.
@@ 4.4 Ablation study
The ablation study in Table 5 evaluates the impact of key components and hyperparameters on model performance. Removing the semantic loss $\mathcal{L}_{\text{sem}}$ degrades text quality, increasing PPL by 23.6% while retaining strong robustness. Disabling the online editor $N$ significantly weakens resistance to text edits, with F1 scores dropping by 5.3% for paraphrasing and 2.7% for the copy-paste attack. Our model enables users to balance quality metrics flexibly by adjusting the top-$k$ logits tokens without retraining. Increasing $k$ from 20 to 40 enhances log diversity but results in a slight increase in perplexity. Adjusting $\delta$ reveals a clear trade-off between robustness and text quality. Lowering $\delta$ to 0.75 reduces F1 scores for paraphrasing to 0.854 but achieves the lowest PPL at 7.057. Conversely, increasing $\delta$ improves robustness, with F1 scores peaking at 0.962 for paraphrasing and 0.992 for the copy-paste attack at $\delta=1.5$, though at the cost of higher PPL, a 21.3% increase compared to $\delta=0.75$. The optimal balance is achieved at $\delta=1.25$, where F1 scores for paraphrasing remain high at 0.916 while maintaining a moderate PPL of 7.730.
@@ 5 Conclusion
We introduce a novel logits-based end-to-end model, where encoder and decoder networks are jointly optimized to improve detection robustness and text quality. Constructing such a system is challenging due to the non-differentiability of key modules, such as the online text editor and semantic loss computation. To overcome these obstacles, we propose a novel online prompting technique that leverages the on-the-fly LLM as a differentiable surrogate, effectively handling these non-differentiable operations. Our method can be easily generalized to different LLMs.
Appendix A More Comparison
Table 6: Performance of the LLM watermark methods. CL: Clean watermark sample; SS: Synonymous substitution; CP: Copy-paste attack; PA: Paragraphing; PPL: text perplexity. The best is in bold, and the second-best is underlined.
Method Llama2-7B NLLB-600M Starcoder
Robustness Quality
CL SS CP PA Avg. PPL BLEU pass@1
@@ Appendix A More Comparison
We compare our method with extra three logits-based methods: SWEET(Lee et al., 2023), SIR(Liu et al., 2024b), and TSW(Huo et al., 2024). Table 6 presents the detection performance and text quality across different watermarking methods on Llama2-7B, NLLB-600M, and Starcoder. The evaluation includes robustness against three editing types (SS, CP, PA) and quality metrics (PPL, BLEU, pass@1). Our method outperforms all competitors, achieving the highest average F1 score of 0.969, surpassing the second-best method, Unigram by 3.9%. It demonstrates exceptional resilience against the copy-paste and paraphrasing attacks, achieving 0.978 and 0.916, respectively, outperforming the strongest baseline (Unigram) in each category by 13.2% and 3.5%. Furthermore, our method achieves the highest BLEU score (31.062), exceeding the next-best approach (Unbiased, 28.949) by 7.3%, and attains a competitive pass@1 score (34.0), outperforming Unigram by 3.0%. In contrast, methods such as KGW and Unigram exhibit strong robustness but at the cost of increased PPL (8.658 and 9.275, respectively), whereas our method maintains a lower PPL (7.730), improving output quality by 16.6% compared to Unigram. These results highlight the effectiveness of our approach in achieving superior watermark robustness while preserving high text quality across diverse tasks.
Appendix B Effect on Sentence Length
Fig. 8 presents the F1 score of various watermarking methods across different sentence lengths for three LLMs: OPT-1.3B, Llama2-7B, and Qwen2.5-7B. The results demonstrate that F1 scores generally improve as sentence length increases, indicating enhanced watermark detectability in longer text. Our method consistently achieves strong performance across all models and sentence lengths. Notably, for shorter sentences (length $\leq 50$), our method attains higher F1 scores compared to other baselines, specially with an average improvement of 8% over DiPmark and Unbiased. For longer sentences (length $\geq 100$), all methods converge toward near-optimal F1 scores ($\approx 1.00$). These findings highlight the robustness of our approach in watermark detection, particularly in challenging short-text scenarios, while maintaining competitive performance across different LLM architectures.
Appendix C Effect on LLM Temperature

### v2

# An End-to-End Model for Logits-Based Large Language Models Watermarking

@@ Abstract
The rise of LLMs has increased concerns over source tracing and copyright protection for AIGC, highlighting the need for advanced detection technologies. Passive detection methods usually face high false positives, while active watermarking techniques using logits or sampling manipulation offer more effective protection. Existing LLM watermarking methods, though effective on unaltered content, suffer significant performance drops when the text is modified and could introduce biases that degrade LLM performance in downstream tasks. These methods fail to achieve an optimal tradeoff between text quality and robustness, particularly due to the lack of end-to-end optimization of the encoder and decoder. In this paper, we introduce a novel end-to-end logits perturbation method for watermarking LLM-generated text. By jointly optimization, our approach achieves a better balance between quality and robustness. To address non-differentiable operations in the end-to-end training pipeline, we introduce an online prompting technique that leverages the on-the-fly LLM as a differentiable surrogate. Our method achieves superior robustness, outperforming distortion-free methods by 37–39% under paraphrasing and 17.2% on average, while maintaining text quality on par with these distortion-free methods in terms of text perplexity and downstream tasks. Our method can be easily generalized to different LLMs. Code is available at https://github.com/KAHIMWONG/E2E_
LLM_WM
@@ 3.3 Online Text Editing
As shown in Fig. 1 (b), the online text editor $N$ (in light purple) is positioned between the encoder and decoder during training to simulate user edits of watermarked content, enhancing detection robustness. We utilize the online prompting technique in Fig. 2 (a) to augment watermarked text by dynamically prompting the online LLM $M$. This approach effectively handles non-differentiable online text modifications, such as rewriting. Specifically, after generates watermarked text $\mathbf{X}_{\text{wm}}$, which is fed into $N$ to produce $\hat{\mathbf{X}}_{\text{wm}}$. The editing prompt: Rewrite the following paragraph: [text] . Now start to rewrite the above paragraph: instructs $M$ to generate an augmented version of the watermarked content, where [text] is a placeholder for $\mathbf{X}_{\text{wm}}$. Thus, $N(\mathbf{X}_{\text{wm}})$ is equivalent to $M([\mathbf{X}_{\text{epb}},\mathbf{X}_{\text{wm}},\mathbf{X}_{\text{epe}}])$ in which $\mathbf{X}_{\text{epb}}$ and $\mathbf{X}_{\text{epe}}$ denote the beginning and the end of the editing prompt. In contrast to existing online text editing methods, such as random token dropping/adding (Zhang et al., 2024), our approach shows significantly improved robustness to unseen distortions due to the capabilities of the on-the-fly $M$ to perform more complex modification on the text (more examples are provided in Appendix C.4). While external paraphrasing models like Dipper (Krishna et al., 2023) are available, differences in tokenizers between the LLM and these models render the operation non-differentiable.
@@ 3.6 Cross-LLM Inference
Robustness (F1↑) Quality Robustness (F1↑) Quality Robustness (F1↑) Quality
@@ 4.1.1 Implementation Details
To train our end-to-end model, we choose OPT-1.3B as the online LLM for efficiency. We use samples from the WikiText-103 dataset (Merity et al., 2017) as prompts for training and C4 (Raffel et al., 2020) for evaluation. The context window is set to be $W=10$, with the watermark strength $\delta=1.25$ and $k=20$ for the top-$k$ watermark logits empirically. We employ MGDA (Huo et al., 2024; Désidéri, 2012) to balance the detection and semantic objectives and use the Adam optimizer with a fixed learning rate of 1e-4 training with 35k steps. All experiments are conducted on one single NVIDIA RTX A6000 48G GPU. The evaluation is performed based on the MarkLLM [^2] https://github.com/THU-BPM/MarkLLM (Pan et al., 2024), an open-source tool for benchmarking LLM watermark methods. For further training and evaluation details, please refer to our code and Appendix C.
@@ 4.1.2 Metrics
Following the tradition of Pan et al. (2024); Zhang et al. (2024), we benchmark LLM watermarking methods across four key aspects: 1) Detection effectiveness of the human-written and clean watermarked text, measured by the F1 scores at optimal thresholds; 2) Detection robustness, evaluated by subjecting the watermarked text to 3 types of text modifications; 3) Text quality, assessed using the text perplexity as well as performance on downstream tasks including machine translation and code generation; and 4) Efficiency of the watermarking model, evaluated by measuring the time and GPU memory overhead for text generation and watermark detection presented in Appendix E.
@@ 4.1.3 Competitors
We compare our method with the SOTA logits-based methods: KGW (Kirchenbauer et al., 2023), SIR (Liu et al., 2024b), and Unigram (Zhao et al., 2024), as well as distortion-free methods: Unbiased (Hu et al., 2024) and DiPmark (Wu et al., 2024). A comparison involve more competitors is presented in Appendix A.1.
@@ 4.2 Robustness Quality Benchmarking
Table 2 compares the robustness and quality of five methods across three LLMs, with Unigram and DiPmark serving as the strong baselines. The performance gap ($\Delta$) is presented at the bottom of the table. We use the first 30 tokens from the C4 dataset (Raffel et al., 2020) as prompts, and generate 200 clean watermarked tokens as a watermark sample, with original human-written text serving as non-watermarked samples. Robustness is evaluated under edited conditions (SS, CP, PA in Table 2), while text quality is assessed using PPL with Llama2-13B as the oracle. Unlike Unbiased, which relies on prompts and LLM API access, making it inefficient, our method and other baselines operate without such constraints. Notably, our model is trained exclusively on OPT-1.3B and employs the converter (see Sec. 3.6) for cross-LLM inference. Our method outperforms all baselines in robustness and quality across most scenarios, with minor degradations (1–3%) in a few cases, achieving an average F1 score of 0.975 across all scenarios. Our method achieves an average robustness improvement of 17.3% over DiPmark, with notable gains of 37–39% under paraphrasing, while maintaining comparable PPL scores. Compared to Unigram, our approach shows a 1.5% average robustness gain, peaking at 12% under copy-paste attacks for Llama2-7B, and reduces PPL by over 9% on average. These results highlight that our method effectively delivers watermark resilience while preserving LLM output quality, making it a practical solution for real-world applications.
@@
To further validate the LLM-agnostic generalizability of our method, we conduct experiments on three additional LLMs, with the results shown in Table 3. Our approach consistently achieves a high F1 score ($\geq 0.99$) on clean watermarked samples across all LLMs, demonstrating its stability and reliability. Moreover, it maintains strong robustness against all editing types, yielding an average F1 score of 0.969. In terms of text quality, our method introduce only a moderate increase in PPL, approximately 1.2× that of non-watermarked baselines. These zero-shot results highlight the effectiveness of our approach without necessitating LLM-specific tuning, underscoring its broad applicability across diverse LLM architectures.
@@
Machine translation task with NLLB-600M
@@
Code generation task with Starcoder
@@
The strength parameter $\delta$ in logits-based methods is a critical factor in balancing robustness and quality. To systematically evaluate this trade-off, we conduct experiments assessing watermark resilience against paraphrasing and the copy-paste attack while measuring PPL. Watermarked text is generated under varying $\delta$ values [^3] Each method is tested at its default strength, as well as one lower and one higher setting: KGW and Unigram: {1, 2, 5}; SIR: {0.5, 1, 2}; Ours: {0.75, 1.25, 1.5}. The results, presented in Fig. 5, demonstrate the superior performance of our method in achieving an optimal balance between robustness and quality. At a PPL threshold of 8, where the non-watermarked baseline has a PPL of 6.8, our approach improves robustness by 18.16% under paraphrasing and 20.42% under copy-paste edits, surpassing the second-best method, Unigarm. In contrast, existing approaches such as KGW, Unigram, and SIR exhibit significant text quality degradation, with PPL values increasing by two to three times compared to the default $\delta$ while offering only marginal robustness improvements. These results highlight the effectiveness of our method in preserving text quality while achieving substantial robustness against text modifications.
Additional quantitative results are presented in Appendix A.
@@ 4.3 Comparsion to KGW-type Methods
We justify our proposed neural-based method by comparing to the KGW-type approaches, highlighting that our method offers guarantees analogous to provable p-values, which are essential for text forensics.
Red-Green List Retrieval. Although our method does not explicitly rely on the p-value to determine the presence of the watermark, we allow for forensic analysis by retrieving the logits perturbation, which serves as the counterpart to the red-green list in KGW, for each token in a given sentence. Specifically, given an input sequence $[\mathbf{x}_{1},\dots,\mathbf{x}_{n}]$, tokenization is performed using the proposed converter. The retrieval is achieved by first obtaining the output distribution of $M$ at each step for $[\mathbf{x}_{w},\dots,\mathbf{x}_{n}]$ and subsequently reconstructing $\mathbf{S}_{\text{top}\,k}^{(t)}$ for each step $t$. The logits perturbation are then sequentially derived as $\mathbf{l}^{(t)}_{\text{top}\,k}=E(\mathbf{S}_{\text{top}\,k}^{(t)})$. Our watermarked samples colored with retrieval logits perturbation are visualized in Appendix B.1, this retrieval process enables the classification of tokens into green and red lists, thereby enhancing the interpretability of our watermark signal.
@@
Statistical Decoder vs. Neural Decoder. Our method enables the retrieval of logits perturbation for a given sentence, allowing us to visualize the relationship between cumulative perturbed logits and neural decoder logits in Fig. 6. The analysis shows a strong positive correlation, where neural decoder logits increase as cumulative perturbed logits rise across the two LLMs. Additionally, the neural decoder outperforms the statistical decoder, likely due to the joint optimization process, which allows it to capture latent features that enhance detection performance. The detection performance of our statistical and neural decoder is presented in the Appendix B.2.
False Positive Thresholding. Our model uses a neural decoder to generate a scalar confidence score for watermark detection. Similar to KGW-type methods, both employ a fixed threshold to detect watermarks, enabling adjustments to control false positive rates. Fig. 7 shows ROC curves under various text modifications, demonstrating the model’s ability to manage false positive rates effectively. FPR across more LLMs is shown in Appendix A.5.
@@ 4.4 Ablation study
The ablation study in Table 5 evaluates the impact of key components and hyperparameters on model performance. Removing the semantic loss $\mathcal{L}_{\text{sem}}$ degrades text quality, increasing PPL by 23.6% while retaining strong robustness. Disabling the online editor $N$ significantly weakens resistance to text edits, with F1 scores dropping by 5.7% for paraphrasing and 2.7% for the copy-paste attack. Our model enables users to balance quality metrics flexibly by adjusting the top-$k$ logits tokens without retraining. Increasing $k$ from 20 to 40 enhances log diversity but results in a slight increase in perplexity. Adjusting $\delta$ reveals a clear trade-off between robustness and text quality. Lowering $\delta$ to 0.75 reduces F1 scores for paraphrasing to 0.854 but achieves the lowest PPL at 7.057. Conversely, increasing $\delta$ improves robustness, with F1 scores peaking at 0.962 for paraphrasing and 0.992 for the copy-paste attack at $\delta=1.5$, though at the cost of higher PPL, a 21.3% increase compared to $\delta=0.75$. The optimal balance is achieved at $\delta=1.25$, where F1 scores for paraphrasing remain high at 0.916 while maintaining a moderate PPL of 7.730.
@@ 5 Conclusion
We introduce a novel logits-based end-to-end model, where encoder and decoder networks are jointly optimized to improve detection robustness and text quality. Challenges stem from the non-differentiability of key modules, the online text editor and semantic loss computation, we overcome by a novel online prompting technique that leverages the on-the-fly LLM as a differentiable surrogate. Our method can be easily generalized to different LLMs.
Acknowledgements
This work was supported in part by Macau Science and Technology Development Fund under SKLIOTSC-2021-2023, 0022/2022/A1, and 0119/2024/RIB2; in part by Research Committee at University of Macau under MYRG-GRG2023-00058-FST-UMDF; in part by the Guangdong Basic and Applied Basic Research Foundation under Grant 2024A1515012536.
Impact Statement
Embedding resilient watermarks in AI-generated text transforms how we govern and trust automated content. Platforms can reliably verify authorship, deterring plagiarism and malicious reuse, while publishers gain confidence in sourcing and provenance. Content consumers benefit from clearer attribution, reducing the spread of misinformation. Regulators and standard-setting bodies acquire practical tools to enforce intellectual-property rights and audit generative systems. By elevating accountability and transparency across the AI ecosystem, the proposed watermark method approach paves the way for responsible innovation, bolsters public trust, and safeguards creative and journalistic integrity.
Appendix
[…]
Appendix A More Quantitative Results
A.1 More Competitors
Table 6: Overall performance of the LLM watermark methods. CL: Clean watermark sample; SS: Synonymous substitution; CP: Copy-paste attack; PA: Paragraphing. Detection Performance is evaluated on Llama2-7B. PPL: Text perplexity with Llama2-13B; MT: Machine translation with NLLB-600M; CG: Code generation with Starcoder. The best is in bold, and the second-best is underlined.
Method F1↑ PPL↓ BLEU↑ pass@1↑
CL SS CP PA Avg. Qlt MT CG
@@ A.1 More Competitors
We compare our method with extra three logits-based methods: SWEET (Lee et al., 2023), SIR (Liu et al., 2024b), and TSW (Huo et al., 2024). Table 6 presents the detection performance and text quality across different watermark methods. In terms of robustness, our method outperforms all competitors, achieving the highest average F1 score of 0.969, surpassing the second-best method, Unigram by 3.9%. Our method demonstrates exceptional resilience against the copy-paste and paraphrasing attacks, achieving 0.978 and 0.916, respectively, outperforming the strongest baseline (Unigram) in each cases by 13.2% and 3.5%. Furthermore, our method achieves the highest BLEU score of 31.062, exceeding the next-best approach (Unbiased, 28.949) by 7.3%, and attains a competitive pass@1 score of 34.0, outperforming Unigram by 3.0%. In contrast, methods (such as KGW and Unigram) exhibit strong robustness generally cost of increased PPL (8.658 and 9.275, respectively), whereas our method maintains a lower PPL pf 7.730, improving output quality by 16.7% compared to Unigram. These results highlight the effectiveness of our approach in achieving superior watermark robustness while preserving high text quality across diverse tasks.
A.2 Effect on Sentence Length